given the point P(-3,5),Q(5,1),S(1,-2),R(5,6)find the coordinates of the mid_point of PQ
Answer by stanbon(75887)

given the point P(-3,5),Q(5,1),S(1,-2),R(5,6)find the coordinates of the mid_point of PQ
------
Let the midpoint be (x,y)
----
x = (-3+5)/2 = 2/2 = 1
y = (5+1)/2 = 6/2 = 3
----
Ans:: (1,3)
------------------
Cheers,
Stan H.
==============